OpenStudy (anonymous):

An airplane over the Pacific Ocean sights an atoll at an 13° angle of depression. If the plane is 752 m above the water, how many kilometers is it from a point 752 m directly above the atoll?

OpenStudy (jhannybean):



OpenStudy (jhannybean):

Can you solve this using \(\large\tan (13)=\frac{752}{x}\) ?

OpenStudy (anonymous):

how do you solve it from here i understand how to set it up!

OpenStudy (jhannybean):

First,do you understand why we're using tan and not Sin or cos?

OpenStudy (anonymous):

yes!

OpenStudy (jhannybean):

ok :D \[\large \tan(13)=\frac{752}{x}\]\[\large \frac{1}{752} \cdot \tan(13) = \frac{752}{x}\cdot \frac{1}{752}\]\[\large x = \frac{\tan(13)}{752}\]

OpenStudy (jhannybean):

Did ya get it? :P

OpenStudy (anonymous):

I think i did it wrong! the answer choices are: A. 3.26 km B. 173.61 km C. 1624.12 km D. 3257.27 km